# Let’s do the time-warp-attend: Learning topological invariants of dynamical systems

### 19 Jan 2024 v2

@@ Abstract
Keywords: dynamical systems, bifurcations, topological invariance, Hopf bifurcation, physics-informed machine learning, augmentation, single-cell RNA-sequencing.
@@ 1 Introduction
Understanding bifurcations, be they in physics, engineering, or biology, is a subject of intense research in the study of nonlinear systems (Kuznetsov, 1998; Karniadakis et al., 2021; Ghadami & Epureanu, 2022). The central challenge across all applications is that bifurcations can be qualitatively identical (e.g. the emergence of oscillations) while being quantitatively very different (e.g. the amplitude or frequency of those oscillations); smoothly warping the state space has no qualitative effect on the dynamics. Put another way, a bifurcation detector must be selective for topological structure of underlying dynamics but invariant to the geometric nuisances of those dynamics. Predicting the onset of these bifurcations often requires the explicit use of governing equations, which must be hardwired or inferred. Hence, the analysis of real data, where such equations are often unknown or difficult to define, is particularly challenging (Kuznetsov, 1998). Applications to real data either use hardwired (Wissel, 1984; Wiesenfeld, 1985; Scheffer et al., 2009) or learned (Bury et al., 2023) time series features, but no approach, to our knowledge, explicitly addresses the selectivity-invariance trade-off inherent to the topological nature of the problem. Additionally, where vector field data is readily available and time-series tracing is not, e.g., in capturing gene expression of cells in high-throughput single-cell experiments, it remains unclear how to extract topological information either from such noisy and sparse vector fields directly or from their integration.
We address this challenge by learning a universal dictionary of dynamical features which is explicitly encouraged to be invariant to geometric nuisances but selective for the topological features which define a particular type of bifurcation (Fig. 1). This is achieved primarily by the use of topological data augmentation: each sample of our training data is a velocity field which comes bundled with numerous deformed but identically labeled versions. The features extracted from these augmented samples are then passed to a supervised classifier which must contend with the substantial geometric nuisances induced by the augmentation scheme. We demonstrate our method for the ubiquitous supercritical Hopf bifurcation whereby a stable fixed point (or “equilibrium”) loses stability and transitions into a regime exhibiting a stable limit cycle. We train on synthetic, augmented data generated from a simple oscillator prototype. Despite being only trained on data from one simple equation, our architecture can generalize to radically different systems, such as the nonlinear Selḱov glycolysis oscillator and the Belousov-Zhabotinsky model of chemical oscillations, precisely because of our use of topological augmentation.
@@ 2 Related Work
Classical methods for detecting limit cycles Solving non-linear differential equations is in general difficult, and it has been recognized since the time of Poincaré that reasoning about dynamical systems’ characteristics can be made easier by qualitatively delineating solutions according to their long-term behaviors. For autonomous (i.e., time-invariant) systems in two dimensions, if the dynamics do not diverge, these behaviors are either fixed points or limit cycles (see Methods), and transitions between these dynamical regimes as a result of changing parameters are called bifurcations. More explicitly, a bifurcation corresponds to a change in the topological structure of a phase portrait as a parameter of the dynamical system is varied. Full characterization of the location and number of limit cycles in two-dimensional, polynomial systems is a challenging open problem in dynamical systems theory (Hilbert, 2019; Smale, 2000).
@@
Early theoretical results guaranteeing the existence or absence of a limit cycle in two-dimensional autonomous systems are the theorems of Poincaré-Bendixson and Bendixson-Dulac, respectively. Further guarantees are possible when the system can be reduced to the form of a so-called Liénard equation. None of these results, however, is constructive or easy to use in practice. Center manifold theory (Carr, 2012), for its part, analyzes dynamics on representative subspaces where systems reduce to normal forms having known bifurcation behaviors, though this method requires knowledge of the underlying governing equation. The same limitation applies to Conley-Morse theory, which represents dynamics as a graph whose nodes correspond to long-term behaviors (Conley, 1978). Importantly, none of these approaches is adapted to real data, where equations are unknown and system measurements can be sparse, incomplete and noisy.
@@
Data-driven methods for detecting limit cycles Data-driven methods fall into non-topological and topological camps. In the former camp, several approaches use compressed sensing of trajectory data to reduce noise, followed by identification of system parameters and associated bifurcation diagrams (Wang et al., 2011; 2016), though this method requires inferring governing equations for each new sample. Other work sought to reduce vector field measurements to a Conley-Morse graph (Chen et al., 2007b; a) for which a database over parameter regimes can be constructed via combinatorial dynamics exploration and graph equivalence comparisons (Arai et al., 2009). Yet other approaches used estimates of Lyapunov exponents to delineate between periodic and stationary systems, although, as we show below, this approach can be sensitive to noise (Stoker, 1950; Zhang et al., 2009).
@@
On the topological front, there is a growing body of work using persistent homology to detect periodicity in time series data (Ravishanker & Chen, 2021). For example, Perea & Harer (2015) extracted persistence barcodes from delay embeddings of time series to measure periodicity, although this process is very computationally expensive and there is no standard approach to classifying barcodes. Persistent homology approaches often assume that the underlying bifurcation parameter is known (Tymochko et al., 2020), which greatly simplifies the problem at the cost of realism. To our knowledge, there is no data-driven method which is efficient (not requiring expensive barcode calculations for each sample), general (applicable to novel systems without retraining) and robust (not sensitive to noise).
@@
Generalization across dynamical systems. Comparatively few approaches exist for the simultaneous analysis of diverse dynamical systems and their bifurcation behavior. Work by Brown & Sethna (2003); Transtrum & Qiu (2014) and Quinn et al. (2021) mapped families of dynamical systems to a manifold whose geometry encodes different qualitative behaviors, but these methods have no straightforward way of identifying bifurcations between these behaviors. More recent work from Ricci et al. (2022) uses a data-driven approach called phase2vec to learn manifolds of dynamical systems, leading to dynamical embeddings which encode qualitative properties of the underlying systems, though it is not clear if this method can detect or distinguish dynamical regimes and their associated bifurcations per se.
@@
Diffeomorphic augmentations for generalization. Diffeomorphic transformations which smoothly preserve topology in the context of dynamical systems have been used principally in computer vision. In order to train a classifier of a limited image set of biological entities such as diatoms, Vallez et al. (2022) used diffeomorphic augmentations to expand its train data, however, since the labels of these data were not necessarily preserved by these transformations, a manual post-filtering was required to maintain realism. In a second set of works concerning dynamical systems specifically, Khadivar et al. (2020) and Bevanda et al. (2022) fit a diffeomorphic transformation into closed or linearized forms, and work by Bramburger et al. (2021) used an autoencoder approach to reduce a complex dynamical system to a topologically conjugate (i.e., dynamically equivalent) form of discrete-time mapping. Here, we use diffeomorphisms to curate a set of consistently labeled dynamical augmentations, see Appendix 7.1.1.
@@ 3 Methodology
Topological equivalence in dynamical systems Let $\dot{x}=f(x;\theta)$ for $x\in\mathbb{R}^{2}$ be a parameterized dynamical system with parameter vector $\theta\in\Theta$. A solution to this equation, $F(x_{0},t)$, with an initial condition, $x(t=0)=x_{0}$, is called a trajectory through $x_{0}$. A standard approach to understanding dynamical systems (Kuznetsov, 1998) is to characterize them by the topology of all of their trajectories, construed as functions of $x$ and $t$ or just $x$. For example, do trajectories extend indefinitely, stop at a point (a fixed point) or loop back on themselves (a limit cycle)? More precisely, we say that $f$ has a fixed point at $x^{*}$ and $\theta=\theta^{*}$ if $f(x^{*};\theta^{*})=0$. We say that $f$ has a limit cycle at $\theta^{*}$ if there exists a trajectory of $f$ which is closed and isolated (neighboring trajectories are not closed). Importantly, these properties are preserved when the state space is subjected to a topology-preserving map, making them topological invariants of the dynamical systems. However, these properties may change as $\theta$ changes; i.e. the system undergoes a bifurcation. For example, in a system that undergoes a supercritical Hopf bifurcation, there exists a parameter $\theta$ such that below a certain threshold of $\theta$ the system exhibits a stable fixed point, and above it, the fixed point loses its stability and a stable limit cycle emerges.
Dynamical equivalence and topological augmentation
In what follows, we will consider autonomous systems with either a single stable fixed point (i.e., equilibrium) or a stable limit cycle with an unstable fixed point in its interior. We refer to systems, $f$, in one of these two categories as belonging to the same dynamical class, denoted $y(f)$, and we denote the fixed point and limit cycle classes considered here the “pre-Hopf” and “post-Hopf” classes, respectively. As discussed above, these classes are invariant to topology-preserving deformations of the state space. Here, we will use diffeomorphisms, $h\in H$ which induce an equivalence relation on dynamical systems in a standard and well-studied way: we have $y(f)=y(f\circ h)$ for all $h\in H$ so that a given $H$ defines an equivalence relation which partitions the space of dynamical systems into its dynamical classes. We refer to two systems, $f,g$ such that $g=f\circ h$ as being dynamically equivalent  [^2] We choose the neutral term “dynamical equivalence” to distinguish our notion from the similar notions of topological conjugacy and topological equivalence (Kuznetsov, 1998). See 7.1.1 for details..
@@
Our training data is constructed by first acquiring planar vector fields, $\{f(x,\theta)\in\mathbb{R}^{2}\}$ evaluated at all points, $x$, on a fixed, $64\times 64$ lattice. Parameters, $\theta$, are sampled uniformly on $a\in[-0.5,0.5],\omega\in[-1,1]$, and each system is assigned a label according to whether it is in the pre- or post-Hopf dynamical class. Then, the dataset is augmented with expressive diffeomorphisms generated using monotonic rational-quadratic splines. In particular, we use the spline construction from Durkan et al. (2019), where instead of training the neural spline flows proposed in that work, we use its random initialization which generates a diverse set of diffeomorphic augmentations. For details on data augmentation, see Appendix 7.1.2.
Convolutional attention training. We construe our full, augmented data set as a collection of labeled arrays in $\mathbb{R}^{2\times 64\times 64}$ which we use to train a convolutional network. Our architecture is adapted from the discriminator of self-attention generative adversarial network (SAGAN) of Zhang et al. (2019), coupled with a multi-layered perception (MLP). For standardization of the data, we reduce each vector field to only its angular components, ignoring magnitudes which carry little information about cyclic behavior. The feature extraction module consists of four convolutional blocks with self-attention mechanism into the last two convolutional layers. Our reasoning for incorporating self-attention is that only a few dynamical keypoints in the vector field (i.e., the fixed point or the limit cycle) carry information about class label, and the rest of the array can be (nearly) ignored. Together, the data warping and attentional components of our pipeline comprise our eponymous method. Features from this network are passed to a multilayer perceptron classifier which is trained with a supervised, cross-entropy loss. Because of our augmentation scheme, this classifier must contend with data which, though radically different in appearance, belong to the same dynamical class, encouraging the classifier to learn the true dynamical invariances underlying the Hopf bifurcation. Training and architecture details are described in the Appendix.
@@ 4.1 Transferring knowledge of dynamical classes from simple prototypes to complex test data
We evaluate our model and its baseline competitors on a diverse, synthetic dataset of noisy (zero-mean gaussian), real-world systems drawn from across the sciences. Test datasets include the simple harmonic oscillator (“SO”), its warped variant (“Augmented SO”), a system undergoing a supercritical Hopf bifurcation as described in (“Supercritical Hopf”), examples based on Liénard equations (“Liénard Polynomial” and “Liénard Sigmoid”), and systems used to model oscillatory phenomena in nature and engineering applications, namely the van der Pol system, used for modeling electrical circuits in early radios (“Van der Pol”), the Belousov-Zhabotinsky model, a model of chemical oscillations (“BZ reaction”), and the Selḱov oscillator, used in modeling glycolitic cycles (“Selḱov”), see Appendix 7.2 and Strogatz (1994).
@@
Our baselines comprise several competing feature types: (1) “Critical Points”, a heuristic based on critical points identification as derived from Helman & Hesselink (1989; 1991) algorithm, (2) “Lyapunov” which distinguishes pre- and post-Hopf states according to the value of the Lyapunov exponent of a time series integrated from the vector field, (3) “Phase2vec”, vector field embeddings from Ricci et al. (2022), a state-of-the-art deep-learning method for dynamical systems embedding, and (4) “Autoencoder”, latent features of of a convolutional vector field autoencoder (the same as that used in phase2vec training), see Appendix 7.3. As with our method, for all baselines except Critical Points, classifications are produced by a linear classifier trained on features from augmented simple harmonic oscillator data (Augmented SO) in order to discriminate between point and cyclic attractors. The accuracy average and standard deviation for our model and the trained baselines are computed over 50 training runs.
@@
We find that our framework generalizes across a wide range of dynamical systems and classifies distinct dynamical regimes in noisy (zero-mean gaussian) conditions better on average than baseline methods (see Table A4, Table 1 and examples at Fig. A8). Critical Points achieves nearly perfect classification on noiseless data but is easily disrupted by noise (e.g., the perfect classification of Supercritical Hopf drops to 56% with Gaussian noise $\sigma=0.1$). Likewise, Lyapunov suffers a substantial decrease in accuracy for classical oscillators (e.g., 89% accuracy of Supercritical Hopf reduces to 70% accuracy with Gaussian noise $\sigma=0.1$). Both with and without noise, Phase2vec extends mainly to SO and Supercritical Hopf. Autoencoder obtains competitive results, however, it exhibits little generalization to Liénard Polynomial system, and essentially no generalization to Selḱov system. We also test resilience to noise in the Augmented SO data (see Table 2). Despite the substantial distortion of the vector fields caused by noise, the accuracies of our model and that of autoencoder representation only gradually decrease, unlike the sharp decrease in accuracy for the other baselines.
@@
In the Appendix 7.3, we report the results of another baseline, “Parameters” which trains a linear classifier based on a flattened vector of coefficients of a degree three polynomial fit to each vector field with least squares, a common method for bifurcation analysis, although this resulted in a poor average accuracy of 53%.

### 21 Mar 2024 v3

@@ Abstract
Keywords: dynamical systems, bifurcations, topological invariance, Hopf bifurcation,
physics-informed machine learning, augmentation, single-cell RNA-sequencing.
@@ 1 Introduction
Understanding bifurcations which transition between dynamical regimes, be they in physics, engineering, or biology, is a subject of intense research in the study of nonlinear systems (Kuznetsov, 1998; Karniadakis et al., 2021; Ghadami & Epureanu, 2022). One of the central challenges across all applications is that distinct dynamical regimes can be qualitatively identical (e.g. the emergence of oscillations) while being quantitatively very different (e.g. the amplitude or frequency of those oscillations); smoothly warping the state space has no qualitative effect on the dynamics. Put another way, a detector for different dynamical regimes and the transitions between them must be selective for topological structure of underlying dynamics but invariant to the geometric nuisances of those dynamics. Predicting the onset of such bifurcations often requires the explicit use of governing equations, which must be hardwired or inferred. Hence, the analysis of real data, where such equations are often unknown or difficult to define, is particularly challenging (Kuznetsov, 1998). Applications to real data either use hardwired (Wissel, 1984; Wiesenfeld, 1985; Scheffer et al., 2009) or learned (Bury et al., 2023) time series features, but no approach, to our knowledge, explicitly addresses the selectivity-invariance trade-off inherent to the topological nature of the problem. Additionally, where vector field data is readily available and time-series tracing is not, e.g., in capturing gene expression of cells in high-throughput single-cell experiments, it remains unclear how to extract topological information either from such noisy and sparse vector fields directly or from their integration.
We address this challenge by learning a universal dictionary of dynamical features which is explicitly encouraged to be invariant to geometric nuisances but selective for the topological features which define a particular type of bifurcation (Fig. 1). This is achieved primarily by the use of topological data augmentation: each sample of our training data is a velocity field which comes bundled with numerous deformed but identically labeled versions. The features extracted from these augmented samples are then passed to a supervised classifier which must contend with the substantial geometric nuisances induced by the augmentation scheme. We demonstrate our method for the ubiquitous supercritical Hopf bifurcation whereby a stable fixed point (or “equilibrium”) loses stability and transitions into a regime exhibiting a stable limit cycle. We train on synthetic, augmented data generated from a simple oscillator prototype. Despite being only trained on data from one simple equation, our architecture can generalize to radically different systems, such as the nonlinear Selḱov glycolysis oscillator and the Belousov-Zhabotinsky model of chemical oscillations.
@@ 2 Related Work
Classical methods for detecting limit cycles. Solving non-linear differential equations is in general difficult, and it has been recognized since the time of Poincaré that reasoning about dynamical systems’ characteristics can be made easier by qualitatively delineating solutions according to their long-term behaviors. For autonomous (i.e., time-invariant) systems in two dimensions, if the dynamics do not diverge, these behaviors are either fixed points or limit cycles, and transitions between these dynamical regimes as a result of changing parameters are called bifurcations. More explicitly, a bifurcation in these cases corresponds to a change in the topological structure of a phase portrait as a parameter of the dynamical system is varied.
@@
Full characterization of the location and number of limit cycles in two-dimensional, polynomial systems is a challenging open problem in dynamical systems theory (Hilbert, 2019; Smale, 2000). Early theoretical results guaranteeing the existence or absence of a limit cycle in two-dimensional autonomous systems are the theorems of Poincaré-Bendixson and Bendixson-Dulac, respectively. Further guarantees are possible, for example, when the system can be reduced to the form of a so-called Liénard equation. None of these results, however, is constructive or easy to use in practice. Center manifold theory (Carr, 2012), for its part, analyzes dynamics on representative subspaces where systems reduce to normal forms having known bifurcation behaviors, though this method requires knowledge of the underlying governing equations. The same limitation applies to Conley-Morse theory, which represents dynamics as a graph whose nodes correspond to long-term behaviors (Conley, 1978). Importantly, none of these approaches is easily adapted to real data, where equations are in general unknown and system measurements can be sparse, incomplete and noisy.
@@
Data-driven methods for detecting limit cycles. Data-driven methods fall into non-topological and topological classes. In the former class, several approaches use compressed sensing of trajectory data to reduce noise, followed by identification of system parameters and associated bifurcation diagrams (Wang et al., 2011; 2016), though this approach requires inferring governing equations for each new sample. Other work sought to reduce vector field measurements to a Conley-Morse graph (Chen et al., 2007b; a) for which a database over parameter regimes can be constructed via combinatorial dynamics exploration and graph equivalence comparisons (Arai et al., 2009). Yet other approaches used estimates of Lyapunov exponents to delineate between periodic and stationary systems, although, as we show below, this approach can be sensitive to noise (Stoker, 1950; Zhang et al., 2009).
@@
On the topological front, there is a growing body of work using persistent homology to detect periodicity in time series data (Ravishanker & Chen, 2021). For example, Perea & Harer (2015) extracted persistence barcodes from delay embeddings of time series data to measure periodicity, although this process is computationally expensive and there is no standard approach to classifying barcodes. Persistent homology approaches often assume that the underlying bifurcation parameter is known (Tymochko et al., 2020), which greatly simplifies the problem at the cost of realism. To our knowledge, there is no data-driven method which is efficient (not requiring expensive barcode calculations for each sample), general (applicable to novel systems without retraining) and robust (not sensitive to noise).
@@
Generalization across dynamical systems. Comparatively few approaches exist for the simultaneous analysis of diverse dynamical systems and their bifurcation behavior. Work by Brown & Sethna (2003); Transtrum & Qiu (2014) and Quinn et al. (2021) mapped families of dynamical systems to a manifold whose geometry encodes different qualitative behaviors, but these methods have no straightforward way of identifying bifurcations between these behaviors. More recent work from Ricci et al. (2022) uses a data-driven approach called phase2vec to learn manifolds of dynamical systems, leading to dynamical embeddings which encode qualitative properties of the underlying systems, though it is not tailored for the detection of distinct dynamical regimes and their associated bifurcations per se.
@@
Diffeomorphic augmentations for generalization. Diffeomorphic transformations which preserve topology in the context of dynamical systems have been used principally in computer vision. In order to train a classifier of a limited image set of biological entities such as diatoms, Vallez et al. (2022) used diffeomorphic augmentations to expand its train data, however, since the labels of these data were not necessarily preserved by these transformations, a manual post-filtering was required. In a second set of works concerning dynamical systems specifically, Khadivar et al. (2020) and Bevanda et al. (2022) fit a diffeomorphic transformation into closed or linearized forms, and work by Bramburger et al. (2021) used an autoencoder approach to reduce a complex dynamical system to a topologically conjugate (i.e., dynamically equivalent) form of discrete-time mapping. Here, we use diffeomorphisms to curate a set of consistently labeled dynamical augmentations.
@@ 3 Methodology
Topological equivalence in dynamical systems. Let $\dot{x}=f(x;\theta)$ for $x\in\mathbb{R}^{2}$ be a parameterized dynamical system with parameter vector $\theta\in\Theta$. A solution to this equation, $F(x_{0},t)$, with an initial condition, $x(t=0)=x_{0}$, is called a trajectory through $x_{0}$. A standard approach to understanding dynamical systems (Kuznetsov, 1998) is to characterize them by the topology of all of their trajectories, construed as functions of $x$ and $t$ or just $x$. For example, do trajectories extend indefinitely, stop at a point (a fixed point) or loop back on themselves (a limit cycle)? More precisely, we say that $f$ has a fixed point at $x^{*}$ and $\theta=\theta^{*}$ if $f(x^{*};\theta^{*})=0$. We say that $f$ has a limit cycle at $\theta^{*}$ if there exists a trajectory of $f$ which is closed and isolated (neighboring trajectories are not closed). Importantly, these properties are preserved when the state space is subjected to a topology-preserving map, making them topological invariants of the dynamical systems. However, these properties may change as $\theta$ changes; i.e. when the system undergoes a bifurcation. For example, in a system that undergoes a supercritical Hopf bifurcation, there exists a parameter $\theta$ such that below a certain threshold of $\theta$ the system exhibits a stable fixed point, and above it, the fixed point loses its stability and a stable limit cycle emerges. See Appendix 7.1 for definitions.
Dynamical equivalence and topological augmentations.
In what follows, we will consider autonomous systems with either a single stable fixed point (i.e., equilibrium) or a stable limit cycle with an unstable fixed point in its interior. We term these two regimes as dynamical classes, denoted $y(f)$, and we denote the fixed point and limit cycle classes considered here the “pre-Hopf” and “post-Hopf” classes, respectively. As discussed above, these classes are invariant to topology-preserving deformations of the state space. Here, we will use diffeomorphisms, $h\in H$ which induce an equivalence relation on dynamical systems in a standard and well-studied way: we have $y(f)=y(f\circ h)$ for all $h\in H$ so that a given $H$ defines an equivalence relation which partitions the space of dynamical systems into its dynamical classes. We refer to two parameterized systems, $f,g$ such that $g=f\circ h$ as being dynamically equivalent  [^2] We choose the neutral term “dynamical equivalence” to distinguish our notion from the similar notions of topological conjugacy and topological equivalence (Kuznetsov, 1998). See Appendix 7.1.2 for details..
@@
Our training data is constructed by first acquiring planar vector fields, $\{f(x,\theta)\in\mathbb{R}^{2}\}$ evaluated at all points, $x$, on a fixed, $64\times 64$ lattice. Parameters, $\theta$, are sampled uniformly on $a\in[-0.5,0.5],\omega\in[-1,1]$, and each system is assigned a label according to whether it is in the pre- or post-Hopf regime. Then, the dataset is augmented with diffeomorphisms generated using monotonic rational-quadratic splines. In particular, we use the spline construction from Durkan et al. (2019), where instead of training the neural spline flows proposed in that work, we use its random initialization which generates a diverse set of diffeomorphic augmentations. For details on data augmentation, see Appendix 7.2.1.
Convolutional attention training. We construe our full, augmented data set as a collection of labeled arrays in $\mathbb{R}^{2\times 64\times 64}$ which we use to train a convolutional network. Our architecture is adapted from the discriminator of self-attention generative adversarial network (SAGAN) of Zhang et al. (2019), coupled with a multi-layered perception (MLP). For standardization of the data, we reduce each vector field to only its angular components, ignoring magnitudes which carry little information about cyclic behavior. The feature extraction module consists of four convolutional blocks with self-attention mechanism into the last two convolutional layers. Our reasoning for incorporating self-attention is that only a few dynamical keypoints in the vector field (i.e., the fixed point or the limit cycle) carry information about class label, and the rest of the array can be (nearly) ignored. Together, the data warping and attentional components of our pipeline comprise our eponymous method. Features from this network are passed to a multilayer perceptron classifier which is trained with a supervised, cross-entropy loss. Because of our augmentation scheme, this classifier must contend with data which, though radically different in appearance, belong to the same dynamical class, encouraging the classifier to learn the true dynamical invariances underlying the Hopf bifurcation. Training and architecture details are described in Appendix 7.2.2.
@@ 4.1 Transferring knowledge of dynamical classes from simple prototypes to complex test data
We evaluate our model and its baseline competitors on a diverse, synthetic dataset of noisy (zero-mean gaussian), real-world systems drawn from across the sciences. Test datasets include the simple harmonic oscillator (“SO”), its warped variant (“Augmented SO”), a system undergoing a supercritical Hopf bifurcation as described in (Strogatz, 1994)(“Supercritical Hopf”), examples based on Liénard equations (“Liénard Polynomial” and “Liénard Sigmoid”), and systems used to model oscillatory phenomena in nature and engineering applications, namely the van der Pol system, used for modeling electrical circuits in early radios (“Van der Pol”), the Belousov-Zhabotinsky model, a model of chemical oscillations (“BZ reaction”), and the Selḱov oscillator, used in modeling glycolitic cycles (“Selḱov”), see Appendix 7.3 and Strogatz (1994).
@@
Our baselines comprise: (1) “Critical Points”, a heuristic based on critical points identification as derived from Helman & Hesselink (1989; 1991) algorithm, (2) “Lyapunov” which distinguishes pre- and post-Hopf classes according to the value of the Lyapunov exponent of a time series integrated from the vector field, (3) “Phase2vec”, vector field embeddings from Ricci et al. (2022), a state-of-the-art deep-learning method for dynamical systems embedding, and (4) “Autoencoder”, latent features of a convolutional vector field autoencoder, see Appendix 7.4. As with our method, for all baselines except Critical Points, classifications are produced by a linear classifier trained on features from augmented simple harmonic oscillator data (Augmented SO) in order to discriminate between point and cyclic attractors. The accuracy average and standard deviation for our model and the trained baselines are computed over 50 training runs.
@@
We find that our framework generalizes across a wide range of dynamical systems and classifies distinct dynamical regimes in noisy (zero-mean gaussian) conditions better on average than baseline methods (see Table 1, Table A4, and examples in Fig. A8). Critical Points achieves nearly perfect classification on noiseless data but is easily disrupted by noise (e.g., the perfect classification of Supercritical Hopf drops to 56% with Gaussian noise $\sigma=0.1$). Likewise, Lyapunov suffers a substantial decrease in accuracy for classical oscillators (e.g., 89% accuracy of Supercritical Hopf reduces to 70% accuracy with Gaussian noise $\sigma=0.1$). Both with and without noise, Phase2vec extends mainly to SO and Supercritical Hopf. Autoencoder obtains competitive results but exhibits little generalization to Liénard Polynomial system, and essentially no generalization to Selḱov system. We also test resilience to noise in the Augmented SO data (see Table 2). Despite the substantial distortion of the vector fields caused by noise, the accuracies of our model and that of autoencoder representation only gradually decrease, unlike the sharp decrease in accuracy for the other baselines.
@@
In Appendix 7.4, we report the results of another baseline, “Parameters” which trains a linear classifier based on a flattened vector of coefficients of a degree three polynomial fit to each vector field with least squares, a common method for bifurcation analysis, although this resulted in a poor average accuracy of 53%.